OpenStudy (anonymous):

I don't understand how to get the point (x,y) on the unit circle that corresponds to the real number t. t=5π/4

OpenStudy (anonymous):

count \[\frac{5\pi}{4}\] around the unit circle. the first coordinate will be \[-\frac{\sqrt{2}}{2}\] as will the second

OpenStudy (anonymous):

http://tutorial.math.lamar.edu/cheat_table.aspx

OpenStudy (anonymous):

look at unit circle on the last page of the trig cheat sheet and you will see them all mapped out

OpenStudy (anonymous):

a unit circle has parametric equation: x=sint, y=cost. this t is the real value called parameter. in ur example it is 5*pi/4. just substitute and get the co-ordinates.

OpenStudy (anonymous):

rather x= cost, y= sint

OpenStudy (anonymous):

i think the point of knowing the points on the unit circle is to be able to find the sine and cosine of \[\frac{5\pi}{4}\] not the other way around

OpenStudy (anonymous):

ya. Maybe u r right........